Oconee Federal Financial Corporation has paid a constant annual dividend payment of $1.70 per share. Suppose that you expect that Oconee Federal Financial Corporation will continue to pay this constant dividend payment for the foreseeable future.

(a) What is the most you would be willing to pay for a share of Virtusa Corporation if you have a required return of 12%?

(b) Suppose that the board of directors of Oconee Federal Financial Corporation makes a surprise announcement that they plan to grow the dividend payment to 2% annually from now on. Given this new information, what is the most you would be willing to pay for a share of Oconee Federal Financial Corporation if you have a required return of 12%?
Answer #1

Present Value of a Perpetuity = Annual Payment ÷ Discount Rate

(A). Share Value = 1.70 / 0.12

= $14.17

(B) Perpetuity with growth formula

PV = C / (r – g)

= 1.7 / (0.12 - 0.02)

= 1.7 / 0.10

= $17